# Parametric Factorization of Matrices

### v1

@@ Abstract.
In this survey paper we study parametric versions of writing a matrix in $SL_{n}(\mathbb{C})$ as a product of lower and upper unitriangular matrices in interchanging order as well as generalizations to other classical groups. We give an account of algebraic, continuous and holomorphic factorization results, from the standpoint of Several Complex Variables. Out of the wealth of algebraic results, we only concentrate on those which are related to holomorphic factorization and often formulate them in a specific form, e.g. for the field of complex numbers in place of more general fields or principal ideal domains. The number of unitriangular matrices needed is a difficult problem and is solved in very specific cases only. We give a new lower bound for factorizing matrices in $SL_{2}(\mathbb{C})$ continuously parametrized by two dimensional normal topological spaces.
@@ 1. Introduction
In Section 2 we present some results which depend on general algebraic or topological properties of the ring $R$. We give examples of polynomial rings or function rings with these properties, thus showing the relevance of the presented results for the question of parametric factorization. On the other hand, for continuous or holomorphic dependence of parameter, many function rings do not satisfy these properties. Instead, one should impose the nullhomotopy condition. We discuss the role of nullhomotopy in parametric factorization for continuous, holomorphic and even algebraic dependence in Section 3. Next, in Section 4 we give specific results over polynomial rings. In Section 5 we collect all known results on rings of continuous complex-valued functions. Then in Section 6 we give known factorizations for rings of holomorphic functions on Stein spaces. When a factorization into unitriangular factors exists, it is natural to ask about the number of factors needed. Finding the optimal number is generally a difficult task. We comment on this for all results from Section 2 to 6. The last section, Section 7, is devoted to a generalization of parametric factorization into another direction, the setting of vector bundle automorphisms. We are not aware of any general formulation of the vector bundle setting in terms of algebraic K-theory. In the last section we propose some future development and open problems.
@@ 2.2. Bass stable rank 1
where $U(A_{1},R)$ and $U^{-}(A_{1},R)$ are the upper and lower unitriangular matrices, respectively.
@@
When $R$ has Bass stable rank one, we conclude that $SL_{n}(R)=E_{n}(R)$ and $Sp_{2n}(R)=Ep_{2n}(R)$ for all $n\geq 2$. The next result shows that the number of factors needed for the factorization does not increase with the size of the matrix.
@@ 5. Continuous results
For any natural number $n$ and an integer $d\geq 0$ there exists a natural number $L=L(n,d)$ such that for any finite dimensional normal topological space $T$ of dimension $d$, every nullhomotopic continuous map $f\colon T\to SL_{n}(\mathbb{C})$ can be written as a product of no more than $K$ unitriangular matrices with entries in $\mathcal{C}(T)$.
@@
From the proof of Theorem 5.3 in Christopher Phillips [MR1305876] and the inequality in Proposition 4.1 of [MR4678654] one gets the following lower bound for the number $L$ in Theorem 5.1.
@@
The following lower bound for the two dimensional case seems to be new.
@@
We give an example where four factors are not sufficient. Before we describe the example, let us give some general consideration on factorization of continuous map into four factors.
Given a continuous map $f\colon X\to SL_{2}(\mathbb{C})$, suppose that there were continuous maps $g_{1},g_{2},g_{3},g_{4}\colon X\to\mathbb{C}$ such that
@@
$T=\{(z,w,t)\in\mathbb{C}^{2}\times[0,1]:zw=2-t,\lvert z\rvert=1\}\cong S^{1}\times[0,1].$
@@
From the proof of Theorem 5.3 in Christopher Phillips [MR1305876] and the inequality in Proposition 4.1 of [MR4678654] (with the same proof for the symplectic case) one gets the following lower bound for the number $L_{sp}$ in Theorem 5.5.
@@ 6. Holomorphic results
Regarding the number of factors, first we have the Bass stable one case $K(n,1)=4$ for all $n\geq 1$, cf. Theorem 2.3 and Theorem 2.9 in Section 2.2. Again Theorem 2.7 implies
@@
From Phillips’ lower bound for continuous factorization into exponential factors [MR1305876], Brudnyi and Amol Sasane recently deduced the following lower bound for unitriangular factorization.
@@
$\left(\begin{matrix}1+zw&z^{2}\\ -w^{2}&1-zw\end{matrix}\right)\begin{pmatrix}1&0\\ \frac{1+zw-e^{zw}}{z^{2}}&1\end{pmatrix}=\begin{pmatrix}e^{zw}&z^{2}\\ *&1-zw\end{pmatrix}$.
@@
Anagolous to the $SL_{n}$ case, Corollary 13.6 in Brudnyi–Sasane [BrudSasa] and Proposition 4.1 of [MR4678654] imply the following lower bound for unitriangular factorization of symplectic matrices.
@@ 8. Future developments, open problems
It is also of interest to have more accurate estimates for the number of factors for the various factoriaztions discussed in this survey.

### v2

@@ Abstract.
In this survey paper we study parametric versions of writing a matrix in $SL_{n}(\mathbb{C})$ as a product of lower and upper unitriangular matrices in interchanging order as well as generalizations to other classical groups. We give an account of algebraic, continuous and holomorphic factorization results, from the standpoint of Several Complex Variables. Out of the wealth of algebraic results, we only concentrate on those which are related to holomorphic factorization and often formulate them in a specific form, e.g. for the field of complex numbers in place of more general fields or principal ideal domains. The number of unitriangular matrices needed is a difficult problem and is solved in very specific cases only. We give a new lower bound for factorizing matrices in $SL_{2}(\mathbb{C})$ continuously parametrized by normal topological spaces of dimension bigger than one.
@@ 1. Introduction
In Section 2 we present some results which depend on general algebraic or topological properties of the ring $R$. We give examples of polynomial rings or function rings with these properties, thus showing the relevance of the presented results for the question of parametric factorization. On the other hand, for continuous or holomorphic dependence of parameter, many function rings do not satisfy these properties. Instead, one should impose the nullhomotopy condition. We discuss the role of nullhomotopy in parametric factorization for continuous, holomorphic and even algebraic dependence in Section 3. Next, in Section 4 we give specific results over polynomial rings. In Section 5 we collect all known results on rings of continuous complex-valued functions. Then in Section 6 we give known factorizations for rings of holomorphic functions on Stein spaces. When a factorization into unitriangular factors exists, it is natural to ask about the number of factors needed. Finding the optimal number is generally a difficult task. We comment on this for all results from Section 2 to 6. Section 7 is devoted to a generalization of parametric factorization into another direction, the setting of vector bundle automorphisms. We are not aware of any general formulation of the vector bundle setting in terms of algebraic K-theory. In the last section we propose some future development and open problems.
@@ 2.2. Bass stable rank 1
where $A_{1}$ denotes the rank one reduced irreducible root system of type $A$ and $U(A_{1},R),U^{-}(A_{1},R)$ are the upper and lower unitriangular matrices, respectively.
@@
When $R$ has Bass stable rank one, we conclude that $SL_{n}(R)=E_{n}(R)$ and $Sp_{2n}(R)=Ep_{2n}(R)$ for all $n\geq 2$.
We recall some terminology from Lie algebra. Let $V$ be a finite dimensional real vector space endowed with a scalar product $(\cdot,\cdot)$. Every nonzero $\alpha\in V$ yields a reflection $w_{\alpha}\colon V\to V,\beta\mapsto\beta-\frac{2(\beta,\alpha)}{(\alpha,\alpha)}\alpha$. A subset $\Phi\subset V$ is called a root system of $V$ if the following conditions are satisfied:
  1. (i)
    $\Phi$ is finite, spans $V$ and does not contain 0.
  2. (ii)
    $\Phi$ is invariant under reflection $w_{\alpha}$ for all $\alpha\in\Phi$.
  3. (iii)
    $2(\beta,\alpha)/(\alpha,\alpha)$ is an integer for all $\alpha,\beta\in\Phi$.
The rank of $\Phi$ is the dimension of $V$. Moreover, a root system $\Phi$ is called
  • reduced, if for each $\alpha\in\Phi$, the only roots proportional to $\alpha$ are $\alpha,-\alpha$.
  • irreducible, if it cannot be decomposed into two nonempty complementary subsets such that each root in one set is orthogonal to each root in the other.
The next result shows that the number of factors needed for the factorization does not increase with the size of the matrix.
@@ 5. Continuous results
For any natural number $n$ and an integer $d\geq 0$ there exists a natural number $L=L(n,d)$ such that for any finite dimensional normal topological space $T$ of dimension $d$, every nullhomotopic continuous map $f\colon T\to SL_{n}(\mathbb{C})$ can be written as a product of no more than $L$ unitriangular matrices with entries in $\mathcal{C}(T)$.
@@
Inspired by the work of Brudnyi and Amol Sasane [BrudSasa], one can deduce a lower bound for $L(n,d)$. In fact using the proof of Theorem 5.3 (replacing elementary by unitriangular) in Christopher Phillips [MR1305876] together with the proof of Lemma 2.1 in Brudnyi [MR4357335], one gets the following lower bound for the number $L$ in Theorem 5.1. Alternatively, one can also directly combine Theorem 2.3 of [MR1305876] and Lemma 2.1 of [MR4357335].
@@
The following lower bound for the (at least) two dimensional case seems to be new.
@@
We give a two dimensional example where four factors are not sufficient. Since $L(n,d+1)\geq L(n,d)$ the claim of the Lemma follows. The latter inequality follows from the fact that if a nullhomotopic continuous map $f\colon T\to SL_{n}(\mathbb{C})$ cannot be written as a product of $L$ unitriangular matrices with entries in $\mathcal{C}(T)$, then the map $F\colon T\times\mathbb{R}\to SL_{n}(\mathbb{C})$ defined by $F(t,x)=f(t)$ cannot be written as a product of $L$ unitriangular matrices with entries in $\mathcal{C}(T\times\mathbb{R})$.
Before we describe the example, let us give some general consideration on factorization of continuous map into four factors. Given a continuous map $f\colon X\to SL_{2}(\mathbb{C})$, suppose that there were continuous maps $g_{1},g_{2},g_{3},g_{4}\colon X\to\mathbb{C}$ such that
@@
$T=\{(z,w)\in\mathbb{C}^{2}:\exists\,\,t\in[0,1]:zw=2-t,\lvert z\rvert=1\}\cong S^{1}\times[0,1].$
@@
As in Theorem 5.2, one gets the following lower bound for $L_{sp}$.
@@ 6. Holomorphic results
Regarding the number of factors, first we have the Bass stable rank one case $K(n,1)=4$ for all $n\geq 1$, cf. Theorem 2.3 and Theorem 2.9 in Section 2.2. Again Theorem 2.7 implies
@@
From Phillips’ lower bound for continuous factorization into exponential factors [MR1305876], Brudnyi and Sasane recently deduced the following lower bound for unitriangular factorization.
@@
$\left(\begin{matrix}1+zw&z^{2}\\ -w^{2}&1-zw\end{matrix}\right)\begin{pmatrix}1&0\\ \frac{e^{zw}-1-zw}{z^{2}}&1\end{pmatrix}=\begin{pmatrix}e^{zw}&z^{2}\\ *&1-zw\end{pmatrix}$.
@@
Analogous to the $SL_{n}$ case, Corollary 13.6 in Brudnyi–Sasane [BrudSasa] and Proposition 4.1 of [MR4678654] imply the following lower bound for unitriangular factorization of symplectic matrices.
@@ 8. Future developments, open problems
It is also of interest to have more accurate estimates for the number of factors for the various factorizations discussed in this survey.